Holtzman Clothiers' stock currently sells for $15 a share. It just paid a dividend of $2.5 a share (i.e., D0 = $2.5). The dividend is expected to grow at a constant rate of 4% a year.

What stock price is expected 1 year from now? Round your answer to two decimal places.
$   

What is the required rate of return? Round your answers to two decimal places. Do not round your intermediate calculations.
%
Answer #1

This question is based on the concept of Constant Growth Dividend Discount Model.

As per Constant Growth Dividend Discount Model the dividends will grow at a constant rate forever.

We will use the formula - Price of stock = D1 / (Re - g)

Where D1 is the dividend for year 1

Price of stock or P0 is the current stock price

Re is required rate of return

g is growth

Calculation of D1

D1 = D0 * (1+growth rate)

D1 = $2.50 * (1+0.04)

D1 = $2.60

Calculation of required rate of return or Re

P0 = D1 / (Re - g)

15 = 2.60 / (Re - 0.04)

15Re – 0.60 = 2.60

15Re = 2.60 + 0.60

15Re = 3.20

Re = 3.20 / 15

Re = .2133 or 21.33%

Therefore required rate of return is 21.33%

Now we need to calculate stock price is expected 1 year from now or P1

P1 = D2 / (Re - g)

We need to calculate D2 or Dividend in year 2

D2 = D1 * (1+growth rate)

D2 = $2.60 * (1+0.04)

D2 = $2.704

Therefore P1 will be

P1 = D2 / (Re - g)

P1 = $2.704 / (0.2133 - 0.04)

P1 = $2.704 / 0.1733

P1 = $15.60

The stock price is expected 1 year from now will be $15.60.